OpenStudy (anonymous):

C has the coordinates of (-3,-2) and D has the coordinates of (1,4). What is the lenght of segment CD?

OpenStudy (anonymous):

The Pythagorean Theorem is very useful here. Find the difference between the two x coordinates and the two y coordinates. Then use the The Pythagorean Theorem: length = squareroot of ((change in x)^2 + (change in y)^2) \[L = \sqrt{(\Delta x)^2+(\Delta y)^2}\]

OpenStudy (anonymous):

... where L is the length of the segment CD. \[\Delta x = 1-(-3) = 4, \Delta y = 4-(-2) = 6\] Then L has to be given \[L = \sqrt{(4^2+6^2)} = 2\sqrt13\]